OpenStudy (anonymous):

find last years salary if after 4% pay raise , this salary is 38,480

OpenStudy (e.cociuba):

Last years salary = L So we have a 4% pay raise...Let R be the Raise. R = 4% times L Last years salary + Raise = New Salary New Salary = 38,480 so 38,480 = L + 4%L 38,480 = L(1 + 4%)

OpenStudy (e.cociuba):

But what is 1 + 4%? We know 100% = 1 and 50% = .5 (one half) so to convert percentage to a fraction is move the decimal place. Since we have 4% we can say 4.0 with the decimal place moved left two spaces which is .04 That 0 is VERY important otherwise you have 40%! So 4% = .04 1+.04 = 1.04 38,480 = 1.04 * L Use your calculator to help you through the rest.

OpenStudy (anonymous):

i love you

OpenStudy (e.cociuba):

Hope it helped. Good luck.